# Low-Latency Video Conferencing System for Geo-Distributed Data Centers

### v1

@@ Abstract
In the face of rising global demand for audio/video meetings, managing traffic across geographically distributed (geo-distributed) data centers presents a significant challenge due to the dynamic and limited nature of inter-DC network performance. Facing these issues, this paper introduces two novel techniques, VCRoute and WMJitter, to optimize the performance of geo-distributed video conferencing systems. VCRoute is a routing method designed for video conferencing data packets. It treats the routing problem as a Multi-Armed Bandit issue, and utilizes a tailored Thompson Sampling algorithm for resolution. Unlike traditional approaches, VCRoute uses predicted end-to-end latency as the routing selection reward for each packet, enabling effective and timely end-to-end latency optimization. In conjunction with VCRoute, we present WMJitter, a watermark-based mechanism for managing network jitter. Leveraging a window-based statistic method, WMJitter enables real-time network jitter estimation, leading to significant reductions in end-to-end delay and an improved balance between latency and loss rate. Evaluations based on real geo-distributed network performance demonstrate the effectiveness and scalability of VCRoute and WMJitter, offering robust solutions for optimizing video conferencing systems in geo-distributed settings.
@@ I Introduction
Due to the global distribution of users, most video conferencing systems are also architected and deployed globally to achieve good scalability and low latency. For example, Zoom has 19 interconnected data centers (DCs) spread across the world and can also scale to public clouds when needed [2]. Existing video conferencing systems may have different architectures. For example, Figure 1 shows two service architectures adopted by popular video conferencing systems [3, 4]. For both architectures, large amounts of data are transferred across geo-distributed data centers, either between endpoints and users (Figure 0(a)) or between multiple endpoints (Figure 0(b)). Our study of the geo-distributed network performance on real cloud show that, the inter-DC network latency can be over 10X higher than that of intra-DC network (see Section II-A). The high variance of inter-DC network performance also cause high jitter, which leads to challenges when reducing the end-to-end latency of geo-distributed video conferencing. We summarize the challenges faced by existing video conferencing systems as follows.
@@
First, geo-distributed overlay networks provide a number of relay paths for video conferencing systems to select the best paths to transmit packets between different service endpoints and user pairs. Existing studies have proposed various offline [5] or adaptive [6] algorithms to find the best routing for video conferencing applications. However, when making the routing decisions, existing methods mainly focus on minimizing the packet transmitting latency from source to target, while in video conferencing applications, the end-to-end latency of a packet contains not only the data transmitting time but also packet reordering time sensitive to network jitter (i.e., Figure 3). Due to the dynamic and heterogeneous network performance in geo-distributed DCs, it is highly possible that the routing decisions which lead to the minimum packet transmitting latency do not guarantee the lowest end-to-end latency. For example, the existing routing method [5] which performs a great deal of exploration to find good routing decisions may cause higher network jitter and thus lead to higher packet reordering latency. To provide low-latency video conferencing in geo-distributed DCs, we need to redesign the packet routing approach in existing video conferencing systems considering the special application features.
@@
First, VCRoute is a packet routing method specifically designed for video conferencing systems. VCRoute treats the routing problem as a Multi-Armed Bandit problem and applies a custom version of Thompson Sampling to solve it. In contrast to the conventional approaches [5], our method leverages the Bayesian framework inherent in Thompson Sampling to derive probability distributions for individual arms. This innovative approach allows us to effectively tackle the highly dynamic nature of geo-distributed network performance. Moreover, we innovate by using predicted end-to-end latency as the reward for the routing selection of each packet. Due to the high-latency of geo-distributed network, it is impractical to wait for the reward of a current packet’s routing choice to be sent back from the receiver’s side. To address this issue, we employ a DNN-based online regression model to accurately predict the end-to-end latency for packets. This ensures that the selection probability for each arm can be updated promptly. Furthermore, to maintain good scalability, we have implemented two strategies to minimize the overhead of VCRoute.
Second, we propose WMJitter, a watermark-based Out-Of-Order (OOP) processing mechanism to handle network jitter. WMJitter leverages the advantages of using watermarks in streaming processing systems, enabling a significant reduction in end-to-end delay. We incorporate a window-based statistic method for real-time network jitter estimation, facilitating timely and accurate adjustments to the watermark. This efficient buffering management provided by WMJitter significantly reduces end-to-end latency, thus enhancing the overall system performance by striking a balance between latency and loss rate. These two methods, VCRoute and WMJitter, offer a comprehensive and scalable solution to address the challenges of performance optimization for video conferencing systems in geo-distributed environments.
We evaluate our techniques on two sets of real-world geo-distributed environments with complementary features. By comparing to state-of-the-art routing and jitter managing approaches, we found that 1) VCRoute greatly outperforms existing approaches when the inter-DC network performance is highly heterogeneous (Section VI-C2); 2) watermark-based jitter management has shown great improvement on average end-to-end latency when the inter-DC network performance is highly dynamic (Section VI-C1). In summary, we make the following contributions:
  • To the best of our knowledge, we are the first to jointly consider packet transmitting latency and jitter during packet routing, which is important to reduce the end-to-end latency for data packets in video conferencing systems.
@@ II-A Geo-Distributed Network Features
Geo-distributed data communications go through the Wide Area Network (WAN), which is much different from intra-DC network due to its high latency and high dynamicity [11]. We have the following observations from measuring inter-DC network performance on the real cloud.
TABLE I: Network RTT (ms) and packet loss rate between Tencent cloud instances located in four geo-distributed DCs.
Shanghai Guangzhou Tokyo Moscow
Shanghai (0,0) (29,0.06%) (90,10.96%) (280,6.53%)
Guangzhou (29,0.07%) (0,0) (79,6.23%) (245,0.68%)
Tokyo (90,10.89%) (78,6.13%) (0,0) (153,0.02%)
Moscow (280,6.38%) (245,0.75%) (154,0.03%) (0,0)
@@
First, the inter-DC network latency is much higher than that of intra-DC network. For example, Table I shows the round trip time (RTT) and packet loss rate measured between Tencent cloud instances located in different geo-distributed DCs [12]. When two DCs are far from each other (e.g., Shanghai and Moscow), the network latency between the DCs can reach several hundreds of milliseconds and hence lead to non-negligible delay to inter-DC data communications.
Second, it is sometimes more preferable to transfer data between DCs through relay paths than direct paths considering data communication latency and quality. For example, according to Table I, the round trip delay between Shanghai and Moscow is 280ms, while it can be reduced to 243ms if the data is first transferred from Shanghai to Tokyo and then from Tokyo to Moscow. Similarly, the packet loss rate of direct data transmission from Shanghai to Moscow is 6.53%, while it can be reduced to 0.70% aggregated if the data is relayed via Guangzhou.
Third, inter-DC network latency can vary greatly over time. We measure the RTT between two instances located in different Tencent cloud DCs for a few hours and plot the variation of RTT along time in Figure 2. The highest RTT between Guangzhou and Tokyo can be over ten times larger than the average. There are also inter-DC links with low performance variations, such as that between Shanghai and Guangzhou. This may depend on the distance between two DCs as well as the traffic on the link.
The dynamicity of network latency in the geo-distributed network can lead to jitter problem, which refers to the variation in the delay of packets as they traverse a network. Jitter can be caused by a variety of factors, including network congestion, packet queuing, and routing issues. High levels of jitter can result in poor network performance and may cause problems for real-time applications such as voice and video conferencing.
@@ II-B Video Conferencing Systems
Many popular video conferencing systems such as Zoom and Webex adopt the single service endpoint architecture shown in Figure 0(a) [3], where a single service endpoint is scheduled for each meeting session and all participants communicate with the same endpoint for video streaming. The endpoint is used to select and forward video/audio streams. Due to the high latency of inter-DC network, this can clearly lead to large delay for users far from the endpoint if without any routing optimizations. As discussed in the last subsection, using other service nodes as relays and providing efficient routing policy between endpoint and users may be able to reduce the delay for packets. However, since data packets in the same stream are routed individually, they may take different paths and encounter varying network latency, hence leading to higher jitter.
Existing network communication systems usually adopt jitter buffers to compensate for the quality degradation caused by jitter. For example, Figure 3 shows the jitter buffer management in webRTC [10], a popular open-source framework for creating video conferencing systems. Data packets travel through the Internet and arrive at users’ end. A jitter buffer at the user’s end temporarily stores incoming packets which are usually out-of-order, and then plays them back by reordering the packets according to their timestamps of generation. Packets arriving too late will be discarded. Jitter buffer absorbs the variability in packet arrival times, allowing the communication system to maintain a consistent quality of service. The buffer size is typically configurable. A larger buffer can provide better jitter compensation, at the expense of increased delay, while a smaller buffer may cause more packets being discarded due to out-of-order. Existing studies [13, 14] have proposed various methods to adapt the buffer size according to network jitter.
@@ II-C Motivations
In a global video conferencing session, data streams are transmitted through WAN composed of multiple autonomous systems. Data packets in a stream are routed individually and thus may end up with different paths. Existing routing methods such as Via [5] mainly focus on selecting the best relay paths for packets in geo-distributed overlay networks [16] to minimize packet transmitting time. However, this does not guarantee low end-to-end latency for packets, due to ignorance of packet reordering lag.
@@
In this paper, we found that it is possible to leverage out-of-order processing (OOP) paradigm to further reduce the latency of video conferencing systems. OOP is an architecture commonly used in streaming processing for flexible and efficient execution of stream queries [19]. Compared to the conventional jitter buffer based approach, OOP allows more packets to come out of the reordering operation in time while ensuring the required order, resulting in a reduction in latency.
@@ III System Overview
We formulate the routing problem as a Multi-Armed Bandit (MAB) problem, where each available path in the geo-distributed environment is an arm of the bandit. Different from existing studies [5], we calculate the reward using the end-to-end latency of a packet rather than the transmitting latency (as motivated by Figure 4). However, due to the high latency of inter-DC network, it could take several hundreds of milliseconds for the accurate reward value to be transferred back, which causes delay to our packet-level routing control. Further, the high dynamicity of inter-DC network latency leads to high variance in the rewards, which can significantly impact the performance and effectiveness of traditional MAB algorithms. To address the above challenges, we carefully adapt the Thompson Sampling algorithm [20] to our MAB problem. Thompson Sampling is a probabilistic algorithm that can effectively handle uncertainty and high variance in the rewards of arms. To overcome the challenge of late feedback of reward, we proposed a DNN-based online regression model to quickly and accurately predict the end-to-end latency of a selected path to feedback the reward in a timely manner. This ensures that the selection probability of each arm can be updated promptly and accurately.
Figure 6: An example of handling out-of-order packets using jitter buffer and watermark mechanisms. Y(x) represents that the arrival of the packet triggers the output of packet x.
Low-latency jitter management. Network jitter determines the degree of disorder of the packets. The purpose of jitter management is to absorb the impact of jitter as much as possible, under a low latency. As mentioned above, existing buffer-based jitter management is essentially doing in-order processing and thus can cause unnecessary delay of packets in the buffer. In this paper, we propose watermark-based jitter management [18] at users’ end. Watermarks are quantitative markers associated with a stream, indicating that no events in the stream will have a timestamp preceding that of the watermark. Thus, we can use watermarks to safely determine when to emit packets and keep track of the progress of packet handling. No extra latency will be introduced due to stragglers.
Figure 6 gives an example to show the differences between the buffer-based and watermark-based jitter management. Consider ten packets arriving at the receiver’s side out-of-order. TS means the generated timestamp of a packet and arrival time indicates when the packet arrives at the receiver. When using jitter buffer to handle out-of-order packets, we start to output packets after the buffer gets full. For example, packet 0 is out at timestamp 30ms when the fourth packet (ID=3) arrives. Packet 2 is the second arrived packet, but since it has to wait for packet 1 to emit first to keep the right order, it waits for 10ms extra in the buffer and emits at timestamp 50ms. Thus, packet 1 can be considered as a straggler to packet 2. On the contrary, if we use watermarks, packet 2 can be immediately output at timestamp 40ms. This is because, when packet 6 arrives at timestamp 40ms, the watermark is updated to 30ms, meaning that no packets will arrive in the future with a TS$<$30ms. Thus, we are safe to output all preceding packets with TS30absent30$\leq$ 30≤ 30ms, namely packet 2, 1 and 3. Packet 2 is no longer blocked by the straggler packet 1. The buffer delay, namely the duration that a packet stays in the buffer, shows that watermarks-based jitter manager is more efficient than jitter buffer. In the following, we introduce how watermarks are maintained and updated for video streams (Section V).
@@ IV VCRoute Scheduler
VCRoute provides a centralized routing scheduler for each packet in the video streams, so that the end-to-end latency of the packets can be minimized. To achieve this goal, VCRoute needs to answer the following questions:
@@ IV-B Making good routing decisions
Given the top-k paths, we design an online routing method to dynamically select the best path for each packet (Stage 3 of Figure 7). As mentioned previously, selecting the best path from the k options can be defined as a classic MAB problem. Each path can be considered as an arm of the bandit, and the end-to-end latency of the path is used as the reward of the arm. The MAB problem has been extensively studied. However, the high-variance and dynamically changing reward distributions of the bandits make it hard if not impossible for traditional MAB algorithms such as UCB1 [21, 5] to converge.
@@ IV-B1 Thompson Sampling algorithm
In this paper, we carefully adapt the Thompson Sampling algorithm (TSA) [20] to solve our MAB problem. TSA uses a Bayesian framework to model the uncertainty in reward distributions of arms. It starts from a guess that the rewards follow a prior distribution. When getting more evidence (real rewards), the TSA iteratively updates the hyperparameters of the prior and use them to maintain a posterior distribution following the same parametric form as the prior distribution. In our problem, we assume that the likelihood of the reward of each arm (i.e., end-to-end latency of each path) follows a Normal distribution with known variance. The variance can be calculated using historical calls collected offline. Then, according to Bayesian probability theory, the prior and posterior probability distributions both follow the Normal distribution with hyperparameters mean $\mu$ and variance $\sigma^{2}$. We use the observed rewards to update the hyperparameters for the posterior distribution.
@@
At each time step, TSA samples reward from the posterior predictive normal distribution of each arm and select the arm with the highest returned value. As a start, we initialize $mu=0$ and $\tau=0.001$ in order to explore every arm at the first stage of the process. TSA transparently combines both exploration and exploitation. Once an arm is tested and a reward is obtained, the belief in the likelihood of the reward of that arm is modified. By employing sampling strategies based on posterior probabilities associated with the optimality of different actions, the algorithm ensures continued exploration of all actions that hold plausible optimality potential while reducing sampling emphasis on actions deemed less likely to be optimal. Throughout the iterative process, the algorithm effectively explores promising actions and incrementally dismisses those believed to have lower performance capabilities.
IV-B2 End-to-end latency prediction
The reward is crucial to the effectiveness of our TSA-based routing algorithm. We proposed a DNN-based online regression model to quickly and accurately predict the end-to-end latency of each selected path to feedback the reward in a timely manner (Stage 4 of Figure 7).
Recall that the end-to-end latency of a packet is composed of two parts, namely the packet transmitting latency and packet reordering latency. Theoretically, the first part is correlated to network latency and the second part is correlated to network jitter. Thus, we build a DNN-based prediction model for end-to-end latency with respect to network latency and jitter. Prior to selecting DNN as our prediction model, we conducted an empirical investigation involving alternative algorithms, such as LSTM [22] and Linear Regression [23]. However, these models are not suitable for our specific problem. Notably, we observed that the end-to-end latency for each packet exhibited relative independence. The impact of transmitting latency and jitter on end-to-end latency was found to be non-linear, with the influence being contingent upon the packet processing at the user’s end (e.g., watermark or jitter buffer). As a result, the LSTM and Linear Regression models failed to yield superior outcomes compared to DNN.
We build a five-layer fully connected DNN model. The input layer is used to encode the input feature. The input feature is a two-dimensional vector $[tl,jt]$, where $tl$ and $jt$ represent the transmitting latency and jitter experienced by a packet, respectively. The output of the model is the predicted end-to-end latency of the given packet. We use three hidden layers to improve the robustness of the prediction model. Each hidden layer has 64 neurons and uses $Relu$ as its activation function.
  • Model training. Denote the training dataset as $(X,Y)$, where $X$ includes a set of $[tl,jt]$ for different (source, destination) pairs and $Y$ includes the returned end-to-end latency of packets. We collect the data from multiple meeting sessions at different periods of time with different sources and destinations. The $tl$ values are measured using the round-trip-time of Ping while the $jt$ is calculated and returned by the receiver using Algorithm 2. The total size of the dataset is 356,186 and we use 80% of the dataset for training and 20% for testing.
  • Model inference. Given a packet, we use the trained DNN model to predict the expected end-to-end latency of the packet. We use the real-time network latency fed back by the destination as the $tl$ input to the prediction model. For the $jt$ data, we use the jitter calculated from the transmit latency of packets sent from the source in the past.
[…]

### v2

# Low-Latency Video Conferencing via Optimized Packet Routing and Reordering

@@ Abstract
In the face of rising global demand for video meetings, managing traffic across geographically distributed (geo-distributed) data centers presents a significant challenge due to the dynamic and limited nature of inter-DC network performance. Facing these issues, this paper introduces two novel techniques, VCRoute and WMJitter, to optimize the performance of geo-distributed video conferencing systems. VCRoute is a routing method designed for audio data packets of video conferences. It treats the routing problem as a Multi-Armed Bandit issue, and utilizes a tailored Thompson Sampling algorithm for resolution. Unlike traditional approaches, VCRoute considers transmitting latency and its variance simultaneously by using Thompson Sampling algorithm, which leads to effective end-to-end latency optimization. In conjunction with VCRoute, we present WMJitter, a watermark-based mechanism for managing network jitter, which can further reduce the end-to-end delay and keep an improved balance between latency and loss rate. Evaluations based on real geo-distributed network performance demonstrate the effectiveness and scalability of VCRoute and WMJitter, offering robust solutions for optimizing video conferencing systems in geo-distributed settings.
@@ I Introduction
Due to the global distribution of users, most video conferencing systems are also architected and deployed globally to achieve good scalability and low latency. For example, Zoom has 19 interconnected data centers (DCs) spread across the world and can also scale to public clouds when needed [2]. Existing video conferencing systems may have different architectures. For example, Figure 1 shows two service architectures adopted by popular video conferencing systems [3, 4]. For both architectures, large amounts of data are transferred across geo-distributed data centers, either between endpoints and users (Figure 1(a)) or between multiple endpoints (Figure 1(b)). The existing studies show that the inter-DC network latency can be up to hundreds milliseconds (see Section II-A). The high variance of inter-DC network performance also cause high jitter, which leads to challenges when reducing the end-to-end latency of geo-distributed video conferencing. Below we summarize the challenges faced by existing video conferencing systems.
@@
First, large cloud providers, with geographically dispersed data centers connected by a managed backbone network, allow the creation of efficient overlay alternatives to the default Internet path [5, 6, 7]. Existing studies have proposed various online [5, 8] algorithms to find the best routing for video conferencing applications. However, when making the routing decisions, existing methods mainly focus on minimizing the packet transmitting latency from source to target, while in video conferencing applications, the end-to-end latency of a packet contains not only the data transmitting time but also packet reordering time sensitive to network jitter (i.e., Figure 3). Due to the dynamic and heterogeneous network performance in geo-distributed DCs, it is highly possible that the routing decisions which lead to the minimum average packet transmitting latency do not guarantee the lowest average end-to-end latency. To provide low end-to-end latency video conferencing, we need to redesign the packet routing approach to find routing path with low and stable transmitting latency.
@@
First, VCRoute is a packet routing method specifically designed for video conferencing systems. VCRoute treats the routing problem as a Multi-Armed Bandit problem and applies a custom version of Thompson Sampling to solve it. In contrast to the previous approaches [5], our method leverages the Bayesian framework inherent in Thompson Sampling to derive probability distributions for individual arms. This innovative approach naturally considers transmitting latency and its variance simultaneously, which allows us to tackle the highly dynamic nature of geo-distributed network performance effectively.
Second, we propose WMJitter, a watermark-based Out-Of-Order (OOP) processing mechanism to handle network jitter. WMJitter leverages the advantages of using watermarks in streaming processing systems, enabling a further reduction in end-to-end delay. Leveraging window-based statistic method for real-time network jitter estimation used in WebRTC NetEQ module [12], we facilitate timely and accurate adjustments to the watermark. This efficient buffering management provided by WMJitter can further reduce end-to-end latency, thus enhancing the overall system performance by striking a balance between latency and loss rate. These two methods, VCRoute and WMJitter, offer a comprehensive and scalable solution to address the challenges of performance optimization for video conferencing systems in geo-distributed environments.
We evaluate our techniques on two sets of real-world geo-distributed environments with complementary features. By comparing to state-of-the-art routing and jitter managing approaches, we found that 1) VCRoute greatly outperforms existing approaches when the inter-DC network performance is highly heterogeneous (Section VI-B2); 2) watermark-based jitter management has shown great improvement on average end-to-end latency when the inter-DC network performance is highly dynamic (Section VI-B1). In summary, we make the following contributions:
  • To the best of our knowledge, we are the first to jointly consider transmitting latency and its variance during packet routing, which is important to reduce the end-to-end latency for data packets in video conferencing systems.
@@ II-A Geo-Distributed Network Features
Geo-distributed data communications go through the Wide Area Network (WAN), which is much different from intra-DC network due to its high latency and high dynamicity [13]. In the following, we present observations from existing studies as well as our measurement of real inter-DC network performance on Tencent cloud, to mimic the deploying environment of a geo-distributed video conferencing system using public network connections.
@@
First, the inter-DC network latency of different regions vary a lot. Existing measurement results of popular public clouds [14, 15, 16] show that the maximum medium value of iter-DC network latency (330ms) can be up to 33 times of the minimum value. And a rich of studies [5, 17, 6, 7] have demonstrate that overlay routing is an applicable and effective way to reduce the iter-DC network latency. Second, inter-DC network latency can vary greatly over time. We measure the RTT between two instances located in different Tencent cloud DCs for a few hours and plot the variation of RTT along time in Figure 2. The highest RTT between Guangzhou and Tokyo can be over ten times larger than the average. There are also inter-DC links with low performance variations, such as that between Shanghai and Guangzhou. This may depend on the distance between two DCs as well as the traffic on the link.
The dynamicity of network latency in the geo-distributed network can lead to jitter problem, which refers to the variation in the delay of packets as they traverse a network. Jitter can be caused by a variety of factors, including network congestion, packet queuing, and routing issues. High levels of jitter can result in poor network performance and may cause problems for real-time applications such as voice and video conferencing. Although packet loss rate between inter-DC network paths can also impact the quality of real-time applications, in this paper, we assume existing techniques such as packet loss concealment [18] can adequately address the issue and focus mainly on the high heterogneity and dynamicity issues of network latency.
@@
Popular video conferencing systems such as Zoom and Webex adopt the single service endpoint architecture shown in Figure 1(a) [3], where a single service endpoint is scheduled for each meeting session and all participants communicate with the endpoint for audio streaming. The endpoint is used to select and forward video/audio streams. Due to the high latency of inter-DC network, this can clearly lead to large delay for users far from the endpoint if without any routing optimizations. Whats ’more, data packets in the same stream using default Internet paths (e.g., BGP-derived) are routed individually, they may take different underlay paths and encounter varying network latency, leading to high jitter. As discussed in the last subsection, using other service nodes as relays and providing efficient routing policy between endpoint and users may be able to reduce the delay and jitter for packets.
Existing network communication systems usually adopt jitter buffers to compensate the quality degradation caused by jitter. For example, Figure 3 shows the jitter buffer management in webRTC [12], a popular open-source framework for creating video conferencing systems. Data packets travel through the Internet and arrive at users’ end. A jitter buffer at the user’s end temporarily stores incoming packets which are usually out-of-order, and then plays them back by reordering the packets according to their timestamps of generation. Packets arriving too late are discarded. Jitter buffer absorbs the variability in packet arrival times, allowing the communication system to maintain a consistent quality of service. The buffer size is an important parameter. A large buffer can provide better jitter compensation, at the expense of increased delay, while a small buffer may cause more packets being discarded due to out-of-order. Existing studies [19, 20] proposed to adapt the buffer size according to network jitter.
Figure 3: Jitter Buffer Manager in webRTC
@@ II-C Motivations
In a global video conferencing session, data streams are transmitted through WAN composed of multiple autonomous systems. Data packets in a stream are routed individually and thus may end up with different underlay paths. Existing routing methods such as Via [5] mainly focus on selecting the best relay paths for packets in geo-distributed overlay networks [21] to minimize packet transmitting time. However, this does not guarantee low end-to-end latency for packets, due to ignorance of packet reordering lag.
@@
In this paper, we found that it is possible to leverage out-of-order processing (OOP) paradigm to further reduce the latency of video conferencing systems. OOP is an architecture commonly used in streaming processing for flexible and efficient execution of stream queries [24]. Compared to conventional jitter buffer based approaches, OOP allows more packets to come out of reordering in time while ensuring the required order, resulting in reduced latency.
@@ III System Overview
We formulate the routing problem as a Multi-Armed Bandit (MAB) problem, where each available path in the geo-distributed environment is an arm of the bandit. Traditional MAB algorithms used in existing studies [5] selects average transmitting latency as the sole optimization objective for routing. However, the high dynamicity of inter-DC network latency leads to high variance in the rewards, which can significantly impact the performance and effectiveness of these algorithms. To address the above challenges, we carefully adapt the Thompson Sampling algorithm [25] to our MAB problem. Thompson Sampling can estimate the distribution function of transmitting latency for each arm, allowing for the simultaneous consideration of both transmitting latency and its variance when routing, which can effectively handle uncertainty and high variance in the rewards of arms.
Low-latency jitter management. Network jitter determines the degree of disorder of the packets. The purpose of jitter management is to absorb the impact of jitter as much as possible, under a low latency. As mentioned above, existing buffer-based jitter management is essentially doing in-order processing and thus can cause unnecessary delay of packets in the buffer. In this paper, we propose watermark-based jitter management [23] at users’ end. Watermarks are quantitative markers associated with a stream, indicating that no events in the stream will have a timestamp preceding that of the watermark. Thus, we can use watermarks to safely determine when to emit packets and keep track of the progress of packet handling. No extra latency will be introduced due to existence of straggler packets.
@@ IV VCRoute Scheduler
VCRoute provides a centralized routing scheduler for each packet in the audio streams, so that the end-to-end latency of the packets can be minimized. To achieve this goal, VCRoute needs to answer the following questions:
@@ IV-B Making good routing decisions
Given the top-k paths, we design an online routing method to dynamically select the best path for each packet (Stage 3 of Figure 6). As mentioned previously, selecting the best path from the k options can be defined as a classic MAB problem. Each path can be considered as an arm of the bandit, and the transmitting latency of the path is used as the reward of the arm. The MAB problem has been extensively studied. However, the high-variance and dynamically changing reward distributions of the bandits could make traditional MAB algorithms such as UCB1 [26, 5] to choose path with low average transmitting latency but high variance.
@@ IV-B1 Thompson Sampling algorithm
In this paper, we carefully adapt the Thompson Sampling algorithm (TSA) [25] to solve our MAB problem. TSA uses a Bayesian framework to model the uncertainty in reward distributions of arms. It starts from a guess that the rewards follow a prior distribution. When getting more evidence (real rewards), the TSA iteratively updates the hyperparameters of the prior and use them to maintain a posterior distribution following the same parametric form as the prior distribution. In our problem, we assume that the likelihood of the reward of each arm (i.e., transmitting latency of each path) follows a Normal distribution with known variance. The variance can be calculated using historical calls collected offline. Then, according to Bayesian probability theory, the prior and posterior probability distributions both follow the Normal distribution with hyperparameters mean $\mu$ and variance $\sigma^{2}$. We use the observed rewards to update the hyperparameters for the posterior distribution.
@@
At each time step, TSA samples reward from the posterior predictive normal distribution of each arm and select the arm with the highest returned value. TSA transparently combines both exploration and exploitation. Once an arm is tested and a reward is obtained, the belief in the likelihood of the reward of that arm is modified. By employing sampling strategies based on posterior probabilities associated with the optimality of different actions, the algorithm ensures continued exploration of all actions that hold plausible optimality potential while reducing sampling emphasis on actions deemed less likely to be optimal.